2. The velocity vector of a particle is v(t) (20+4.0t) i- 15jm/s What is the vector in component form and its speed at t-5s? 3. A golf ball is hit off a tee at the edge of a cliff has position vector r(t) = 18.0t i + (4.00t-4.90t*) j m What is the speed of the ball at t 3.00 s? 4. Extra Credit: Find the velocity vector and speed at t 5 of a particle whose acceleration components are as follows 1 234 5 1 2 3 4 5

4. Rewrite the following vectors in polar form a. 2

plz write out solution I have a hard time understanding typed out solutions thanks
Answer #1

NOTE :  

(1) A Vector can be represented in three forms :

(i) Graphical form

Example : -

(2,3) 0

(ii) Component form : overrightarrow{r} = x , hat{i} + y , hat{j}

Example: overrightarrow{r} = 2 , hat{i} + 2 , hat{j}

(iii) Polar form

In this form, we write the vector by writing its magnitude and the angle made by the vector with positive x - axis in anticlockwise direction .

The vector r can be written as ; (r,e

Where, r= is magnitude of the vector r and

θ = tan-1 (1) is the angle made by the vector with positive x -axis in anticlockwise direction

The x - component will be  x = r cos heta

And the y - component will be y = r sin heta

For the example considered above, r = sqrt{2^{2} + 2^{2}} = 2sqrt{2} and θ 45。

So, = (2V2. 45°)

(B) If the two points on a straight line is known then the equation of the straight line can be written as

2-91 (y-y1) = 92 T2T1

Where, (x1 , y1) and (x2 , y2 ) are the two points on the straight line.  

QUESTIONS :

(2) Given,

V(t) (20+4.0t)i 15j m/s

Now, at t = 5 s :

V(t) (20+4.0 x 5)i 15j m/s

or,   V(t)-40i - 15j m/s

This is the component form of the given vector at t= 5 seconds. In which the x - component of the vector is 40 m/s (i.e. velocity/speed of the particle in x direction ) and the y - component is -15 m/s (the negative sign is indicating that the velocity is in negative y direction). This vector can also be written as

V(t) = (40,-15)

We know that, the magnitude of the velocity vector gives the speed of the particle.  

So, Speed |V(t)l- V(40)2+(-15)21600+225 42.72 m/s

This is the speed of the particle at t = 5 seconds.

(3) Given, the position vector of the golf ball is

r(t) = 18.0t i + (4.00t-4.90t*)j m

We know that the velocity is given by

  V(t) - dt

So, V(t)(18.0(4.00t 4.902)) 18.0(4.00-9.80t)j m/s dt

At t = 3 seconds

  V(t) =18.0 , hat{i} + (4.00 - 9.80 imes 3) hat{j} , , m/s = 18.0 , hat{i} -25.4 , hat{j} , , m/s

So,  Speed = , left | V(t) ight | = , sqrt{(18.0)^{2} + (-25.4)^{2}} = , sqrt{324 + 645.16}

  = , sqrt{969.16} , = , 31.1 , , m/s

This is the speed of the golf ball at t = 3 seconds.

(4)  According to the given plots,

In ax Vs t plot (i.e. the plot of x component of acceleration), the equation of the line can be written as

a_{x} - 2 = rac{4-2}{5-0} , imes (t - 0)   

or, a_{x} - 2 = rac{2}{5} , imes t

or,   a_{x} = rac{2t}{5} , + 2

And, we know that

V_{x} = int a_{x} . dt

or,  V_{x} = int left ( rac{2t}{5} +2 ight ) . dt

or, V_{x} = rac{t^{2}}{5} + 2t

Similarly, from the plot of ay Vs t (i.e. the plot of y component of acceleration) , we can write the equation of line as

a,-3 m/s- (as the line is parallel to the t - axis)

So, V_{y} = int 3 , , dt = , 3t , , m/s

So, we can write velocity vector as

V(t) = V_{x} hat{i} + V_{y} hat{j} = left ( rac{t^{2}}{5} + 2t ight ) hat{i} + 3t , hat{j}

Now, at t = 5 seconds

  V(t) = left ( rac{5^{2}}{5} + 2 imes 5 ight ) hat{i} + 3 imes 5 , hat{j} = 15 , hat{i} + 15 , hat{j} in m/s

So,  SpeedVt)- V152 + 152-15V2 m/s

Questions in the second page :

(4) Vectors in the polar form :

(a) Given vector is overrightarrow{r} = 2 , hat{j}

For this vector,  r = left | overrightarrow{r} ight | = 2

and   heta = tan^{-1} rac{y}{x} = tan^{-1}rac{2}{0} = tan^{-1}(infty ) = 90^{circ}

So,   overrightarrow{r} = (2 , , , 90^{circ})

(b) Given vector is overrightarrow{r} = 5 , hat{i} + 5 , hat{j}

For this vector, r = sqrt{5^{2} + 5^{2}} = 5sqrt{2}

and heta = tan^{-1} , rac{5}{5} = tan^{-1}(1) = 45^{circ}

So,   (5V2, 45%

For any doubt please comment and please give an up vote. Thank you.